OpenStudy (anonymous):

medal, fan and fan testimonial. Solve step by step. (x+y)^2 = x^2 + 2xy + y^2 x=3 y=2

OpenStudy (irishboy123):

what do you want to do? \((x+y)^2 = (x+y)(x+y) = x^2 + 2xy + y^2\)

OpenStudy (anonymous):

first insert 3 for x and 2 for y right

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

the LHS and RHS of you equation are equal. what is it you are trying to do?

OpenStudy (madhu.mukherjee.946):

see (x+y)^2=(3+2)^2=25 and x^2+2xy+y^2=9+12+4=25

OpenStudy (madhu.mukherjee.946):

i hope it helps you